OpenStudy (anonymous):

Find the vertex of the parabola x – 4y^2 + 16y – 19 = 0.

OpenStudy (anonymous):

vertex (h, k)=(2, 3) h=-(b/2a)=16/8=2 k=f(2)=3

OpenStudy (mertsj):

This parabola opens to the right, not up so its vertex is (3,2) not (2,3)